OpenStudy (anonymous):

Use the Rational Zeros Theorem to write a list of all possible rational zeros of the function. f(x) = -2x4 + 4x3 + 3x2 + 18

OpenStudy (alyssa_xo):

what are the factors of the leading coefficient? what are the factors of the constant term?

OpenStudy (anonymous):

-2 and 18

OpenStudy (alyssa_xo):

factors of 2: 1,2 factors of 18: 1,2,3,4,6,18 please note im using this as a reference http://www.sparknotes.com/math/algebra2/polynomials/section4.rhtml http://www.cliffsnotes.com/math/algebra/algebra-ii/polynomial-functions/rational-zero-theorem

OpenStudy (alyssa_xo):

so some examples include 1/1,1/2,2/1,2/2,3/1,3/2,...

OpenStudy (anonymous):

ok thank you! Ill get right to learning more about it :D

OpenStudy (alyssa_xo):

with plus minuses apparently. \(\pm \dfrac{1}1{},\pm \dfrac{1}{2},\pm \dfrac{2}{1},\pm \dfrac{2}{2},\pm \dfrac{3}{1},\pm \dfrac{3}{2},\pm \dfrac{4}{1},\pm \dfrac{4}{2},\pm \dfrac{6}{1},\pm \dfrac{6}{2},\pm \dfrac{18}{1},\pm \dfrac{18}{2},\)

OpenStudy (anonymous):

so the answer is all of those?? (the ones right above) @alyssa_xo